# The global attractor of the inelastic linear Boltzmann equation in a gravity field for Maxwell molecules

### v1

@@ 5.2 Long time behaviour of the moments
We will now show that $x_{3}>x_{1}$, assuming that $x_{1}<x_{\infty}$, where $x_{\infty}$ is the abscissa of the equilibrium point $(x_{\infty},y_{\infty})$ of the system (5.11). Such a statement is equivalent to prove that:
@@
The first step to prove the result is to determine more explicitly the coordinates of the equilibrium point. By definition of this equilibrium, we have:
[…]
@@
By assumption, we considered an initial datum of the ODE system that belongs to the region $R_{4}$, so that we have $x_{1}<x_{\infty}$. Observing that the number $\frac{\mu^{2}}{(\mu+1)^{2}}-1$ is negative, we see that the function $x\mapsto x^{\frac{\mu^{2}}{(\mu+1)^{2}}-1}$ is decreasing on $\mathbb{R}_{+}^{*}$, providing that:
@@
Computing now the right hand side of the last inequality we obtain:
@@
The power of $b/c$ in the previous inequality is:
@@
Therefore, we have proved that, for $x_{1}<x_{\infty}$, we have:
@@ 6 Conclusion
Appendix A $\sigma$-representation of the post-collisional velocities in the linear case
To obtain the equation satisfied by the Fourier transform $\mathcal{F}$ of a solution $f$ of (1.11), it is more convenient to write the collision operator using the so-called $\sigma$-representation.
First, let us recall that the $\sigma$-representation in the case of two colliding particles with respective pre-collisional velocities $v,v_{*}$, and respective post-collisional velocities $v^{\prime},v_{*}^{\prime}$ writes
$\left\{\begin{array}[]{ccc}v^{\prime}&=&\displaystyle{\frac{v+v_{*}}{2}+(1-r)\frac{v-v_{*}}{4}+(1+r)\frac{|v-v_{*}|}{4}\sigma},\\ v^{\prime}_{*}&=&\displaystyle{\frac{v+v_{*}}{2}-(1-r)\frac{v-v_{*}}{4}-(1+r)\frac{|v-v_{*}|}{4}\sigma}.\end{array}\right.$ (A.3)
Moreover we recall that the idea behind the $\sigma$-representation is to make explicit the conservation of momentum during collisions, so that the post-collisional velocities have the form $v^{\prime}=\frac{v+v_{*}}{2}+w$, $v_{*}^{\prime}=\frac{v+v_{*}}{2}-w$, and then, to choose directly the direction of $w$ as the parameter $\sigma$, the norm of $w$ being given by $|v-v_{*}|/2=|v^{\prime}-v_{*}^{\prime}|/2$ due to the conservation of the kinetic energy. In general the decomposition $v^{\prime}=\frac{v+v_{*}}{2}+w=\frac{v^{\prime}+v_{*}^{\prime}}{2}$ constraints $w$ to be equal to $\frac{v^{\prime}-v_{*}^{\prime}}{2}$, which we can express using the scattering relation in the $\omega$-representation. For example in the elastic case, we have
$v^{\prime}-v_{*}^{\prime}=v-v_{*}-2((v-v_{*})\cdot\omega)\omega,$ (A.4)
which is another way to see that the norm of $v^{\prime}-v_{*}^{\prime}$ is the same as the norm of $v-v_{*}$. It is also clear that one can choose an arbitrary direction for $v^{\prime}-v_{*}^{\prime}$ up to choosing appropriately $\omega$. In other words, this last expression gives the connection between the $\sigma$- and the $\omega$-representations, since it gives
$\sigma=\frac{v-v_{*}}{|v-v_{*}|}-2\left(\frac{v-v_{*}}{|v-v_{*}|}\cdot\omega\right)\omega.$ (A.5)
In the inelastic case, the norms of the pre- and post-collisional relative velocities are not equal in general and so the link between $\sigma$ and $\omega$ is less direct. Nevertheless, the approach using the scattering relation enables us to obtain the same conclusion in this case as well. In fact, decomposing $w$ in order to have an orthogonal symmetry appear, we have get
$w=\frac{v^{\prime}-v_{*}^{\prime}}{2}$ $=\frac{v-v_{*}}{2}-\frac{(1+r)((v-v_{*})\cdot\omega)\omega}{2}$
$=\frac{v-v_{*}}{2}+\frac{-2(1+r)(v-v_{*})\cdot\omega\omega}{4}$
$=\frac{v-v_{*}}{2}-\frac{(1+r)(v-v_{*})}{4}+\frac{(1+r)(v-v_{*})-2(1+r)((v-v_{*})\cdot\omega)\omega}{4}$
$=(1-r)\frac{v-v_{*}}{4}+(1+r)\frac{|v-v_{*}|}{4}\sigma.$ (A.6)
In the case of the Lorentz gas, we only keep track of the velocity of the tagged particle moving in a background of fixed scatterers, hence the collision rule involves only one velocity $v$. In such a case, decomposing to have an orthogonal symmetry to appear again, we obtain
$v^{\prime}$ $=v-(1+r)(v\cdot\omega)\omega$
$=v-(1+r)\frac{v}{2}+(1+r)\frac{\left(v-2(v\cdot\omega)\omega\right)}{2}$
$=(1-r)\frac{v}{2}+(1+r)\frac{|v|}{2}\sigma.$ (A.7)
To conclude, we notice that, in the elastic case, the $\sigma$-representation for the scattering relation of a Lorentz gas is simply
$v^{\prime}=|v|\sigma.$ (A.8)
Appendix B Evolution equation for the Fourier transform
Using the results of Appendix A we are now in position to write the equation satisfied by the Fourier transform $\mathcal{F}(t,\xi)=\int_{\mathbb{R}^{d}}f(t,v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v$. Without loss of generality we will here assume that $\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)\omega=1$, which will allow us to simplify he computations. We start from the evolution equation (1.11) of $f$, and analyze the terms one by one. Since we have
$\partial_{t}\mathcal{F}(t,\xi)=\int_{\mathbb{R}^{d}}\partial_{t}f(t,v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v,$ (B.1)
we get
$\partial_{t}\mathcal{F}(t,\xi)=-\int_{\mathbb{R}^{d}}a\cdot\partial_{v}f(t,v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v+\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}\frac{1}{r}b\left(\left|\frac{v}{|v|}\cdot\omega\right|\right)f(t,^{\prime}v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v-\mathcal{F}(t,\xi).$ (B.2)
@@ Appendix B Evolution equation for the Fourier transform
$-\int_{\mathbb{R}^{d}}a\cdot\partial_{v}f(t,v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v=a\cdot\int_{\mathbb{R}^{d}}f(t,v)\partial_{v}e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v=-i(a\cdot\xi)\mathcal{F}(t,\xi).$ (B.3)
As far as the gain term is concerned, we start with applying the change of variable ${}^{\prime}v\to v$ for $\omega$ fixed. This yields
$\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}\frac{1}{r}b\left(\left|\frac{v}{|v|}\cdot\omega\right|\right)f(t,^{\prime}v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v=\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}b\left(\left|\frac{v}{|v|}\cdot\omega\right|\right)f(t,v)e^{-iv^{\prime}\cdot\xi}\mathop{}\!\mathrm{d}v.$ (B.4)
Then, we replace the $\omega$-representation by the $\sigma$-representation. We introduce the collision kernel in $\sigma$-representation $b_{\sigma}$, defined as:
$b_{\sigma}\left(\left|\frac{v}{|v|}\cdot\sigma\right|\right)=b\left(\left|\frac{v}{|v|}\cdot\omega\right|\right)$ (B.5)
@@
$v^{\prime}=v^{\prime}(\sigma)=(1-r)\frac{v}{2}+(1+r)\frac{|v|}{2}\sigma.$ (B.6)
To obtain a closed equation on the Fourier transform $\mathcal{F}$, we use the inversion formula that reads
$f(t,v)=\int_{\mathbb{R}^{d}}\frac{\mathcal{F}(t,w)}{(2\pi)^{d}}e^{iv\cdot w}\mathop{}\!\mathrm{d}w,$ (B.7)
from which the second term becomes
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{v}{|v|}\cdot\sigma\right|\right)\frac{\mathcal{F}(t,w)}{(2\pi)^{d}}e^{iv\cdot w}e^{-i\left[\frac{1-r}{2}v+\frac{1+r}{2}|v|\sigma\right]\cdot\xi}\mathop{}\!\mathrm{d}\sigma\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}v.$ (B.8)
In the last integral, for $v$ and $w$ fixed, we introduce the orthogonal symmetry $R=R_{v,\xi}$, which maps $\frac{v}{|v|}$ to $\frac{\xi}{|\xi|}$, and we consider the change of variables $\sigma=R(\widetilde{\sigma})$. Using the fac that $R$ is an involution, we get that $R$ also sends $\frac{\xi}{|\xi|}$ to $\frac{v}{|v|}$. Hence we have that
@@
$=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\widetilde{\sigma}\right|\right)\frac{\mathcal{F}(t,w)}{(2\pi)^{d}}e^{iv\cdot w}e^{-i\left[\frac{1-r}{2}v\cdot\xi+\frac{1+r}{2}|\xi|\widetilde{\sigma}\cdot v\right]}\mathop{}\!\mathrm{d}\widetilde{\sigma}\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}v.$ (B.9)
Rearranging the terms, the Fourier transform of the gain term can be rewritten as:
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\frac{\mathcal{F}(t,w)}{(2\pi)^{d}}e^{-iv\cdot\left[-w+\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|\sigma\right]}\mathop{}\!\mathrm{d}\sigma\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}v.$ (B.10)
@@
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\varphi(y)e^{-ix\cdot y}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}x=\varphi(0),$ (B.11)
which has to be understood in the sense
$\int_{\mathbb{R}^{d}}\widehat{\varphi}(x)\mathop{}\!\mathrm{d}x=\int_{\mathbb{R}^{d}}\widehat{\varphi}(x)e^{i0\cdot x}\mathop{}\!\mathrm{d}x=(2\pi)^{d}\varphi(0),$ (B.12)
where $\widehat{\varphi}$ is the Fourier transform of $\varphi$. Applying (B.11) to the function
$\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\frac{\mathcal{F}(t,-w+\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|\sigma)}{(2\pi)^{d}}\mathop{}\!\mathrm{d}\sigma,$ (B.13)
@@
$\partial_{t}\mathcal{F}(t,\xi)=-i(a\cdot\xi)\mathcal{F}(t,\xi)+\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\mathcal{F}(t,\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|\sigma)\mathop{}\!\mathrm{d}\sigma-\mathcal{F}(t,\xi),$ (B.14)
@@
$\partial_{t}\mathcal{F}(t,\xi)+\left(1+i(a\cdot\xi)\right)\mathcal{F}(t,\xi)=Q^{+}(\mathcal{F})(t,\xi),$ (B.15)
@@
$\overline{\xi}=\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|\sigma.$ (B.16)
@@
$\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\mathcal{F}(t,\overline{\xi})\mathop{}\!\mathrm{d}\sigma$ (B.17)
To conclude this section, let us observe that only very mild assumptions on the angular collision kernel $b$ are necessary to obtain that the collision integral ${\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\mathcal{F}(t,\overline{\xi})\mathop{}\!\mathrm{d}\sigma}$ is continuous given $\mathcal{F}(t,\cdot)\in\mathcal{C}([0,+\infty),\mathcal{C}_{0}(\mathbb{R}^{d}))$ In fact, if we assume that $\mathcal{F}$ is a continuous function, given $\xi,\xi_{*}\in\mathbb{R}^{d}$ we have that
$\left|\mathcal{F}(\xi)-\mathcal{F}(\xi_{*})\right|=\left|\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|e_{1}\cdot\sigma\right|\right)\left[\mathcal{F}(\overline{\xi}_{s})-\mathcal{F}(\overline{\xi_{*}}_{c})\right]\mathop{}\!\mathrm{d}\sigma\right|,$ (B.18)
@@
$\overline{\xi}_{R}=\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|R_{\xi}(\sigma),\hskip 14.22636pt\text{and}\hskip 14.22636pt(\overline{\xi_{*}})_{R}=\frac{1-r}{2}\xi_{*}+\frac{1+r}{2}|\xi_{*}|R_{\xi_{*}}(\sigma),$ (B.19)
with $R_{\xi}$ (respectively $R_{\xi_{*}}$) the orthogonal symmetry that sends $e_{1}$ to $\xi/|\xi|$ (respectively to $\xi_{*}/|\xi_{*}|$). The conclusion follows from the fact that
$\left|\left(\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|R_{\xi}(\sigma)\right)-\left(\frac{1-r}{2}\xi_{*}+\frac{1+r}{2}|\xi_{*}|R_{\xi_{*}}(\sigma)\right)\right|\rightarrow 0$ (B.20)
@@
Acknowledgements. A. Nota and T. Dolmaire gratefully acknowledges the support by the project PRIN 2022 (Research Projects of National Relevance) - Project code 202277WX43.

### v2

@@ 5.2 Long time behaviour of the moments
We will now show that $x_{3}>x_{1}$, assuming that $x_{1}<x_{\infty}$, where $x_{\infty}$ is the abscissa of the equilibrium point $(x_{\infty},y_{\infty})$ of the system (5.8). Such a statement is equivalent to prove that
@@
The first step to prove the result is to determine more explicitly the coordinates of the equilibrium point. By definition of this equilibrium, we have
$y_{\infty}=bx_{\infty}^{-1/\mu}\hskip 8.53581pt\text{and}\hskip 8.53581pty_{\infty}=cx_{\infty}^{1/(\mu+1)},$ (5.28)
@@
By assumption, we considered an initial datum of the ODE system that belongs to the region $R_{4}$, so that we have $x_{1}<x_{\infty}$. Observing that the number $\frac{\mu^{2}}{(\mu+1)^{2}}-1$ is negative, we see that the function $x\mapsto x^{\frac{\mu^{2}}{(\mu+1)^{2}}-1}$ is decreasing on $\mathbb{R}_{+}^{*}$, providing that
@@
Computing now the right hand side of the last inequality we obtain
@@
The power of $b/c$ in the previous inequality is
@@
Therefore, we have proved that, for $x_{1}<x_{\infty}$, we have
@@ 6 Conclusion
Appendix A Evolution equation for the Fourier transform
Here, we write the equation satisfied by the Fourier transform $\mathcal{F}(t,\xi)=\int_{\mathbb{R}^{d}}f(t,v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v$. Without loss of generality we will here assume that $\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)\mathop{}\!\mathrm{d}\omega=1$, which will allow us to simplify the computations. We start from the evolution equation (1.9) of $f$, and analyze the terms one by one. We have
$\partial_{t}\mathcal{F}(t,\xi)=-\int_{\mathbb{R}^{d}}a\cdot\partial_{v}f(t,v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v+\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}\frac{1}{r}b\left(\left|\frac{v}{|v|}\cdot\omega\right|\right)f(t,^{\prime}\hskip-2.84526ptv)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v-\mathcal{F}(t,\xi).$ (A.1)
@@ Appendix A Evolution equation for the Fourier transform
$-\int_{\mathbb{R}^{d}}a\cdot\partial_{v}f(t,v)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v=a\cdot\int_{\mathbb{R}^{d}}f(t,v)\partial_{v}e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v=-i(a\cdot\xi)\mathcal{F}(t,\xi).$ (A.2)
As far as the gain term is concerned, we start with applying the change of variable $^{\prime}\hskip-1.42262ptv\to v$ for $\omega$ fixed. This yields
$\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}\frac{1}{r}b\left(\left|\frac{v}{|v|}\cdot\omega\right|\right)f(t,^{\prime}\hskip-2.84526ptv)e^{-iv\cdot\xi}\mathop{}\!\mathrm{d}v=\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}b\left(\left|\frac{v}{|v|}\cdot\omega\right|\right)f(t,v)e^{-iv^{\prime}\cdot\xi}\mathop{}\!\mathrm{d}v.$ (A.3)
Then, we replace the $\omega$-representation by the $\sigma$-representation. We introduce the collision kernel in $\sigma$-representation $b_{\sigma}$, defined as
$b_{\sigma}\left(\left|\frac{v}{|v|}\cdot\sigma\right|\right)=b\left(\left|\frac{v}{|v|}\cdot\omega\right|\right)$ (A.4)
@@
$v^{\prime}=v^{\prime}(\sigma)=(1-r)\frac{v}{2}+(1+r)\frac{|v|}{2}\sigma.$ (A.5)
To obtain a closed equation on the Fourier transform $\mathcal{F}$, we use the inversion formula from which the second term becomes
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{v}{|v|}\cdot\sigma\right|\right)\frac{\mathcal{F}(t,w)}{(2\pi)^{d}}e^{iv\cdot w}e^{-i\left[\frac{1-r}{2}v+\frac{1+r}{2}|v|\sigma\right]\cdot\xi}\mathop{}\!\mathrm{d}\sigma\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}v.$ (A.6)
[…]
@@
$=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\widetilde{\sigma}\right|\right)\frac{\mathcal{F}(t,w)}{(2\pi)^{d}}e^{iv\cdot w}e^{-i\left[\frac{1-r}{2}v\cdot\xi+\frac{1+r}{2}|\xi|\widetilde{\sigma}\cdot v\right]}\mathop{}\!\mathrm{d}\widetilde{\sigma}\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}v.$ (A.7)
Rearranging the terms, the Fourier transform of the gain term can be rewritten as
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\frac{\mathcal{F}(t,w)}{(2\pi)^{d}}e^{-iv\cdot\left[-w+\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|\sigma\right]}\mathop{}\!\mathrm{d}\sigma\mathop{}\!\mathrm{d}w\mathop{}\!\mathrm{d}v.$ (A.8)
@@
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\varphi(y)e^{-ix\cdot y}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}x=\varphi(0),$ (A.9)
which has to be understood in the sense ${\int_{\mathbb{R}^{d}}\widehat{\varphi}(x)\mathop{}\!\mathrm{d}x=\int_{\mathbb{R}^{d}}\widehat{\varphi}(x)e^{i0\cdot x}\mathop{}\!\mathrm{d}x=(2\pi)^{d}\varphi(0)}$, where $\widehat{\varphi}$ is the Fourier transform of $\varphi$. Applying (A.9) to the function
$\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\frac{\mathcal{F}(t,-w+\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|\sigma)}{(2\pi)^{d}}\mathop{}\!\mathrm{d}\sigma,$ (A.10)
@@
$\partial_{t}\mathcal{F}(t,\xi)=-i(a\cdot\xi)\mathcal{F}(t,\xi)+\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\mathcal{F}(t,\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|\sigma)\mathop{}\!\mathrm{d}\sigma-\mathcal{F}(t,\xi),$ (A.11)
@@
$\partial_{t}\mathcal{F}(t,\xi)+\left(1+i(a\cdot\xi)\right)\mathcal{F}(t,\xi)=Q^{+}(\mathcal{F})(t,\xi),$ (A.12)
@@
$\overline{\xi}=\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|\sigma.$ (A.13)
@@
$Q^{+}(\mathcal{F})=\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\mathcal{F}(t,\overline{\xi})\mathop{}\!\mathrm{d}\sigma$ (A.14)
We remark that only very mild assumptions on the collision kernel $b$ are necessary to obtain that ${\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|\frac{\xi}{|\xi|}\cdot\sigma\right|\right)\mathcal{F}(t,\overline{\xi})\mathop{}\!\mathrm{d}\sigma}$ is continuous provided that $\mathcal{F}(t,\cdot)\in\mathcal{C}([0,+\infty),\mathcal{C}_{0}(\mathbb{R}^{d}))$ . In fact, assuming that $\mathcal{F}$ is continuous, given $\xi,\xi_{*}\in\mathbb{R}^{d}$ we have that
$\left|\mathcal{F}(\xi)-\mathcal{F}(\xi_{*})\right|=\left|\int_{\mathbb{S}^{d-1}}b_{\sigma}\left(\left|e_{1}\cdot\sigma\right|\right)\left[\mathcal{F}(\overline{\xi}_{s})-\mathcal{F}(\overline{\xi_{*}}_{c})\right]\mathop{}\!\mathrm{d}\sigma\right|,$ (A.15)
@@
$\overline{\xi}_{R}=\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|R_{\xi}(\sigma),\hskip 14.22636pt\text{and}\hskip 14.22636pt(\overline{\xi_{*}})_{R}=\frac{1-r}{2}\xi_{*}+\frac{1+r}{2}|\xi_{*}|R_{\xi_{*}}(\sigma),$ (A.16)
with $R_{\xi}$ (respectively $R_{\xi_{*}}$) the orthogonal symmetry that sends $e_{1}$ to $\xi/|\xi|$ (respectively to $\xi_{*}/|\xi_{*}|$). The continuity follows from the fact that
$\left|\left(\frac{1-r}{2}\xi+\frac{1+r}{2}|\xi|R_{\xi}(\sigma)\right)-\left(\frac{1-r}{2}\xi_{*}+\frac{1+r}{2}|\xi_{*}|R_{\xi_{*}}(\sigma)\right)\right|\longrightarrow 0$ (A.17)
@@
Acknowledgements. A. Nota and T. Dolmaire gratefully acknowledge the support by the project PRIN 2022 (Research Projects of National Relevance) - Project code 202277WX43.